Required information The Foundational 15 [LO4-1, LO4-2, LO4-3, LO4-4, LO4-5] [The following information applies to the questiRequired: 1. Prepare the journal entries to record the raw materials used in production and the direct labor cost incurred. (2. Prepare the journal entry to record the overhead cost applied to production. (If no entry is required for a transaction/evThe June 1 work in process inventory consisted of 4,200 units with $14,440 in materials cost and $12,560 in conversion cost.Foundational 4-3 3. How many units were completed and transferred to finished goods during the period? Completed and transferFoundational 4-5 5. Compute the equivalent units of production for conversion. Equivalent units of productionFoundational 4-6 6. What is the cost of beginning work in process inventory plus the cost added during the period for materiaFoundational 4-7 7. What is the cost of beginning work in process inventory plus the cost added during the period for conversFoundational 4-8 8. What is the cost per equivalent unit for materials? (Round your answer to 2 decimal places.) Cost per equ
Answer #1

Journal entry

No General Journal Debit Credit
a Work in process 153250
Raw material 153250
(To record raw material used)
b Work in process 97500
Direct labor 97500
(To record direct labor)

Journal entry

No General Journal Debit Credit
a Work in process 115000
Manufacturing overhead 115000
(To record overhead)

3) Unit completed and transferred out = 4200+36700-9200 = 31700 Units

4) Equivalent unit of material = 31700+9200 = 40900

5) Equivalent unit of conversion = 31700+(9200*50%) = 36300

6) Cost of beginning WIP and cost added for material = 14440+153250 = 167690

7) Cost of beginning WIP and cost added for conversion = 12560+97500+115000 = 225060

8) Cost per equivalent unit of material = 167690/40900 = 4.1